27) How much pure acid should be mixed with 2 gallons of a 50% acid solution in order to get an 80% acid solution?
Answer #1

Suppose x gallons of pure (100%) solution is mixed

with 2 gallons of a 50% acid solution

Given : final solution is 80% pure

Thus,

Total volume of final solution = x+2

For acid total quantity should be equal in initial solutions and final solutions

Thus,

x*100% + 2*50% = (x+2)*80%

10x + 10 = 8x + 16

2x = 6

x = 3 gallons

So, 3 gallons of pure solution should be mixed